Simplify the radical. Assume that all variables represent nonnegative real numbers.
squre root 169m^21n^8

Answer by Alan3354(69443)

Simplify the radical. Assume that all variables represent nonnegative real numbers.
squre root 169m^21n^8
-----------
169 = 13^2
m^21 = (m^10)^2*m
m^8 = (n^4)^2